OpenStudy (anonymous):

how do i solve this equation? (3(x)^1/2)/2-3/(2(x)^1/2)

OpenStudy (anonymous):

all that equal to 0

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

(3/2)(\sqrt(X)-3/2(\sqrt(x)\

OpenStudy (campbell_st):

if it's equal to zero, it makes it easy add the 2nd term to both sides of the equation \[\frac{3x^{\frac{1}{2}}}{2} = \frac{3}{2x^\frac{1}{2}}\] cross multiply and remember, multiplying the same base, add the powers \[3 \times 2 \times x^{\frac{1}{2} + \frac{1}{2}} = 2 \times 3\] which becomes 6x = 6 hope it helps

OpenStudy (anonymous):

i cant understan step e

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

makes perfect sense now :D